OpenStudy (anonymous):

What is 50% of 700? 50 350 650 750 I will medal and fan.

OpenStudy (anonymous):

just divide 700 by 2

OpenStudy (anonymous):

350

OpenStudy (anonymous):

700 divided by 2 = 350 so you are correct

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

there is a same queston just a sec

OpenStudy (anonymous):

What is 30% of 600? 30 180 300 630

OpenStudy (anonymous):

this one is easy and you dont have to do much work. so what is half of 600?

OpenStudy (anonymous):

om 300??

OpenStudy (anonymous):

yes so you can cross out 300 and 630 as the answer. and you can cross out 30 because you can only have 30% =30 if 600 was 100. So whats the remaining answer?

OpenStudy (anonymous):

180

OpenStudy (anonymous):

yes so thats the answer

OpenStudy (mathstudent55):

To find a percentage of a number, change the percent to a decimal and multiply by the number. To change a percent to a decimal, move the decimal point two places to the left.

OpenStudy (anonymous):

thx Ethanwestbrooks

OpenStudy (mathstudent55):

Examples of percents and decimals: % decimal 100 1 75 0.75 60 0.60 35 0.35 10 0.1 2 0.02

OpenStudy (anonymous):

welcome

OpenStudy (mathstudent55):

Examples: What is 25% of 100? Change 25% to a decimal: 25% = 0.25 Multiply 0.25 * 100 = 25 25% of 100 = 25 What is 38% of 800? Change 38% to a decimal: 38% = 0.38 Multiply: 0.38 * 800 = 304 38% of 800 = 304

OpenStudy (mathstudent55):

Now do the same with your problems: 50% of 700 = 0.50 * 700 = 350 30% of 600 = 0.30 * 600 = 180

OpenStudy (anonymous):

ok mathstudent55 What is 5% of 120? 6 12 60 125

OpenStudy (mathstudent55):

Ok. Follow my explanation above and try to do it on your own. Let me know what you get.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

I can guide you through it. First, change 5% to a decimal. Move the decimal point 2 places to the left. What is 5% as a decimal?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

60

OpenStudy (mathstudent55):